Shift Summary for Eve 7/7/99

S. Elizabeth Turner (eturner@kudzu)
Thu, 8 Jul 1999 00:07:18 -0500


Shift Summary for evening shift, 07/07/99

On Shift: E. Turner, P. Toale

Hours of beam available at all intensities: 6
Hours of beam available at nominal intensity: 6
Hours of E832 data (special runs not included): 5.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Linac problems,TeV studies,
low BETA QPM problems, changing tapes
========================================================================
Runs ended this shift:
13834 BEAM_832
13835 BEAM_832

Tapes brought to FCC this shift: None
========================================================================
Accesses during shift: none
========================================================================
Overview of shift:
Mostly good 832 datataking. Only problem was a slightly high (~1.5)
BOA for much of shift.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13834 15:30 Inheirit run from day shift. No beam.
13834 526 16:00 Beam returns.
13834 18:10 Tapes full, run ends. We change tapes.

13835 18:30 Init.
13835 18:55 DN notes new gas bottles XX.141
13835 20:20 No beam -- scheduled TeV studies. MCR
estimates 1/2 hour.
13835 120 21:17 Beam returns.
13835 21:25 MCR is working on BOA, we should (and do)
see some improvement.
13835 22:56 No beam -- low beta QPM quench, estimate
beam will return by 23:45
23:45 Julie arrives with food -- I am very happy.
13835 23:55 No estimate for beam. We stop the run and
switch to muon mode.
========================================================================
Follow up Notes on Previous Shift: none
========================================================================
Items to Follow up on Next Shift:
Muon run when beam returns.